### v2

# On the original Ulam’s problem and its quantization

Abstract.
We show that under general resonance the classical piecewise linear Fermi-Ulam accelerator behaves substantially different from its quantization, in the sense that the classical accelerator exhibits typical recurrence and non-escaping while the quantum version enjoys quadratic energy growth in general. We also describe a procedure to locate the escaping orbits, though exceptionally rare in the infinite-volume phase space, for the classical accelerators, which in particular include Ulam’s very original proposal and the linearly escaping orbits therein in the existing literature, and hence provide a complete (modulo a null set) answer to Ulam’s original question. For the quantum accelerators, we reveal under resonance the direct and explicit connection between the energy growth and the shape of the quasi-energy spectrum.
Chern Institute of Mathematics and LPMC, Nankai University, Tianjin, China. Email: dongchg@nankai.edu.cn. School of Sciences, Great Bay University, Dongguan, China. Email: zhoujing@gbu.edu.cn.
MSC classes: 37N05, 81Q50 (Primary), 35Q41, 37E10 (Secondary).
Keyword: Fermi acceleration, Fermi-Ulam model, resonance, escaping orbits, quasi-energy spectrum.
1. Introduction
In 1949 Fermi [Fermi49] proposed an acceleration mechanism, as to explain for the existence of high energy particles in cosmic rays, that charged particles get repeatedly reflected by moving magnetic mirrors, arguing that the chances of head-on interactions outnumber the head-tail ones so that the particles would on average get accelerated. Later in 1961 Ulam [Ulam61] extracted the effective mechanism in Fermi acceleration, i.e. a particle bounces between two infinitely heavy walls, one fixed and the other moving periodically (c.f. Fig.1). Ulam [Ulam61] performed numerical simulations with a specific piecewise linear wall motion (i.e. $A=1/\sqrt{2},B=\sqrt{2},T=1$ in Fig.1) and then conjectured that such a model should produce escaping orbits, i.e. those whose energy grows to infinity in time.
Since then, numerous efforts, theoretically and numerically, have been made by physicists and mathematicians in the study of the classical Fermi-Ulam accelerators and we refer to [Dol-FA, GRKT12, LiLi80] and references therein for detailed survey on the subject. The application of KAM theory leads to non-acceleration results for sufficiently smooth Fermi-Ulam accelerators and their variations [Dol08potential, kuOr20, LaLe91, Pu83, Pu94, Zhar2000]. On the other hand, escaping orbits appear when singularities are allowed [deSD12, Zha98, Zhou21], ambient potentials are introduced [AZ15, DS09, Pu77], and also when higher-dimensions are considered [GT08nonauto, GT08ham, KZ23, STR10, Zhou20]. Classical techniques such as Birkhoff normal forms and hyperbolic dynamics play fundamental roles in the study of classical Fermi-Ulam accelerators.
For the quantum Fermi-Ulam accelerators, the parallel non-acceleration results were obtained for generic sufficiently smooth wall motions [karn94, Seba90], with the help of quantum KAM theory [how89, how98]. Quantum chaos and acceleration behavior were expected both in numerics [glr14, joco86] and in theory [DoeRi69, Seba90] when singularities are allowed in the models. The quantum Fermi-Ulam accelerators are one of the earliest examples of quantum systems with time-dependent boundary conditions [DoeRi69], and have then inspired the development of many important techniques in the study of mesoscopic physics [Bas88, cbpwc96], quantum control [lbmw03, MaNa10], and nano-devices [AKM14, PiDBMS15], as well as experimental realizations such as cold atoms in dynamic traps [HalGMDHPN10].
Notably among the existing literature on the subject, the closest related to our setting are a special case of piecewise linear Fermi-Ulam accelerator [Zha98] and its quantization [Seba90]. In a specific piecewise linear classical Fermi-Ulam accelerator with the same particular choice of system parameters as Ulam [Ulam61] (i.e. $A=1/\sqrt{2},B=\sqrt{2},T=1$ in Fig.1), Zharnitsky [Zha98] found explicitly a family of linearly escaping orbits. However, these escaping orbits constitute a null set, leaving unanswered the behavior of the majority of orbits on the phase space which has infinite volume. Besides, it was unclear therein [Zha98] if any other choices of parameters can also create acceleration. This issue of “mystery of lucky parameters” also appears in Šeba’s study of the piecewise linear quantum accelerators where he obtained quantum acceleration with the choice of parameters corresponding to the special $1:1$ resonance case in our definition (c.f. Definition LABEL:def:quantumresonanace and Example LABEL:example:11resonance).
In this paper we study a general class of piecewise linear Fermi-Ulam accelerators. We show that under general resonance conditions on the system parameters, the classical accelerators exhibit typical recurrence and the escaping orbits constitute a zero-measure set on the phase space of infinite volume. Meanwhile, the quantum accelerators under general resonance enjoy quadratic energy growth in general, if not always, and the quasi-energy spectra are absolutely continuous with finitely many components. Our results cover and greatly generalize those in [Zha98] and [Seba90] as special cases.
Also, we provide an explicit procedure on how to locate the escaping orbits, though exceptionally rare, in the classical Fermi-Ulam accelerators under general resonance. In particular, this includes Ulam’s very original proposal [Ulam61] (which is the special resonance case in our Definition 2.5), and the linearly escaping orbits found by Zharnitsky therein [Zha98] correspond to an accelerating segment on a very particular rational invariant curve in our language (c.f. Example LABEL:example:escbdd), and hence we provide a complete answer (modulo a null set) to Ulam’s original question. In addition, for the quantum accelerators, we reveal the direct connection between the energy growth behavior and the shape of the quasi-energy spectra with explicit formulas.
From a technical perspective, in the classical accelerators, the rigidity of the piecewise linearity forces the dynamics to be singular and parabolic, making it impossible to apply classical tools such as KAM theory or hyperbolicity as has done in many of the existing literature. For quantum accelerators, the time dependence, especially the delta kicks in wall motions, prevents direct application of a lot machinery for handling Schrödinger equations, just as in the case of the (in)famous kicked rotators. We combine several techniques from adiabatic theory, infinite ergodic theory, and quantum Floquet theory to tackle the challenges.
The paper is organized as follows. We dissect the presentation of results on the classical piecewise linear Fermi-Ulam models in §2 and the corresponding quantum models in §LABEL:sec:QFU, and conclude with discussions on open problems in §LABEL:sec:conc. In §2, we begin with the adiabatic change of coordinates and the adiabatic normals, which serve as the Poincaré maps for the classical accelerators, and then we prove the typical recurrence results under general resonance. In §LABEL:sec:QFU, we first transfer the time dependence in boundary conditions to time-dependent potentials in the Schrödinger equations, then we derive the Floquet wave propagators, and finally we prove the quadratic energy growth and the absolute continuity of the quasi-energy spectra under general resonance.
2. The Classical Fermi-Ulam models
We consider the original Fermi-Ulam model [Ulam61], where a point particle bounces elastically between two infinitely heavy walls, one is fixed and the other moves periodically (c.f. Fig.1). Let $l(t)$ denote the distance between two walls. In Ulam’s original proposal [Ulam61], the periodic wall motion $l(t)$ is piecewise linear:
$l(t)=\begin{cases}B-kt&t\in[0,T)\\ A+k(t-T)&t\in[T,2T)\end{cases}$
where $A\neq B$ are positive parameters and $k:=\dfrac{B-A}{T}$ is the slope.
Throughout the paper, we assume without loss of generality that $B>A$ and hence $k>0$.
$l(t)$
[…]
Figure 1. The original piecewise linear Fermi-Ulam model
The goal of this section is to show in Theorem LABEL:thm:rec that under general resonance (as in Definition 2.5) the classical accelerators exhibit typical recurrence and the escaping orbits constitute a zero-measure set on the phase space of infinite volume. We also provide an explicit procedure for how to locate these exceptionally rare escaping orbits in the classical Fermi-Ulam accelerators under general resonance. In particular, this includes Ulam’s very original proposal (i.e. $A=1/\sqrt{2},B=\sqrt{2},T=1$) [Ulam61] as a particular case of the special resonance case, and the linearly escaping orbits found by Zharnitsky therein [Zha98] correspond to an accelerating segment on a very particular rational invariant curve (c.f. Example LABEL:example:escbdd).
The vital change of coordinates for the special parameters (i.e. $A=1/\sqrt{2},B=\sqrt{2},T=1$) as in [Zha98] cannot be easily generalized, in the authors’ attempt, to other choices of parameters satisfying general resonance. Instead, we begin, in §2.1, with the adiabatic change of coordinates (c.f. Lemma 2.1) and the adiabatic normals (c.f. Proposition 2.2), which describe the behavior of the particle in roughly one period. Then in §2.2 we show that the adiabatic normal forms possess invariant circles (c.f. Lemma 2.4) and then we reveal the hidden skew product structure (c.f. Proposition LABEL:prop:skewprod) on each invariant circle due to general resonance. Finally in §LABEL:ssec:rec we prove Theorem LABEL:thm:rec of the typical recurrence results for resonant accelerators and also describe an explicit procedure to locate escaping orbits on the rational invariant circles.
2.1. The Adiabatic Normal Forms
In this section the main goal is to derive the adiabatic normal forms (c.f. Proposition 2.2) which describe the momentum change of the particle in roughly one period $\Delta t=2T$ for large energies.
We proceed in two stages: firstly we show, in adiabatic coordinates, that the momentum of the particle remains constant along collisions away from singularities (i.e. Lemma 2.1), and secondly we show, in adiabatic normals, that the momentum changes dramatically when the particle encounters singularities (i.e. Proposition 2.2).
The techniques from adiabatic theory have been successfully applied in the study of various Fermi-Ulam accelerators [deSD12, KZ23, Zhou20].
2.1.1. The Adiabatic Coordinates
We start with the derivation of the adiabatic coordinates and show that the momentum (in adiabatic coordinates) of the particle remains constant along collisions away from singularities.
Let $t$ denote the moment of collision at the moving wall and $v$ the velocity of the particle immediately after collision at the moving wall. We denote by $f$ the collision map that sends one collision $(t_{0},v_{0})$ to the next one $(t_{1},v_{1})$. Then for $v_{0}$ large, we have the following equations of motion between two consecutive collisions:
(1) $\begin{cases}v_{0}(t_{1}-t_{0})=l(t_{0})+l(t_{1})\\ v_{1}=v_{0}-2\dot{l}(t_{1})\end{cases}$
In spirit of the adiabatic theory, we observe that
$v_{1}-v_{0}\approx-2\dot{l}(t_{0}),\;t_{1}-t_{0}\approx\frac{2l(t_{0})}{v_{0}},$
which leads to the following ODE
$\frac{dv}{dt}=\frac{-v\dot{l}}{l},$
suggesting that $I:=lv$ is almost constant. Then we update the algorithm by replacing $v$ with $I=lv$ and look for the next-order adiabatic invariant. On the other hand, the new “time” $\theta$ needs to behave in the following way
$\theta_{1}-\theta_{0}=\dfrac{C}{I},$
which leads to another ODE (we aim for constant $C=1$ for simplicity)
${\theta}^{\prime}\frac{2l}{v}=\frac{1}{lv},$
[…]
Ideally, the Euler scheme above would terminate in finite steps since we are dealing with (piecewise) linear wall motions $l(t)$. We propose the following formulas for the adiabatic coordinates and verify in Lemma 2.1 that this is indeed true.
We introduce the new (half) period $\mathcal{T}:=\int_{0}^{T}\dfrac{ds}{l(s)^{2}}=\dfrac{T}{AB}$. We define the adiabatic coordinates $(\theta,I)$ as follows
(2) $\begin{cases}I(t,v):=\mathcal{T}(l(t)v+l(t)\dot{l}(t))\\ \theta(t):=\dfrac{1}{2\mathcal{T}}\displaystyle\int_{0}^{t}\dfrac{ds}{l(s)^{2}}.\end{cases}$
Now we verify that $(\theta,I)$ as defined in (2) are indeed adiabatic in the sense that the new momentum $I$ is constant along collisions away from singularities.
Let $S_{0}$ $(S_{T})$ denote the singular collisions occurring at $t=0$ ($t=T$):
$S_{0}:=\{(t,v)|t=0\},\;S_{T}:=\{(t,v)|t=T\}.$
Let $R_{0}$ ($R_{T}$) be the singularity strip bounded by $S_{0}$ ($S_{T}$) and $f(S_{0})$ ($f(S_{T})$) in the $(t,v)$-phase cylinder. $R_{0}$ ($R_{T}$) is, in fact, the Poincaré section that collects the very first collisions right after the moment $t=0$ ($t=T$) (c.f. Fig.2). The name “adiabatic” manifests itself in the following lemma.
Lemma 2.1 (Adiabatic coordinates).
For $(t_{0},v_{0})\notin f^{-1}(R_{0})\cup f^{-1}(R_{T})$ and $v_{0}$ sufficiently large, the collision map $f$ in adiabatic coordinates $(\theta,I)$ as defined in (2) is given by
(3) $\theta_{1}=\theta_{0}+\frac{1}{I_{0}},\;I_{1}=I_{0}.$
Proof.
We present the proof of the case when $t_{0},t_{1}\in(0,T)$; the other case when $t_{0},t_{1}\in(T,2T)$ can be obtained in a similar fashion.
First in the momentum direction
$\mathcal{T}^{-1}(I_{1}-I_{0})$ $=(l_{1}v_{1}+l_{1}\dot{l}_{1})-(l_{0}v_{0}+l_{0}\dot{l}_{0})$
$=(l_{1}v_{1}-l_{0}v_{0})+(l_{1}\dot{l}_{1}-l_{0}\dot{l}_{0})$
$=((l_{0}-k(t_{1}-t_{0}))(v_{0}+2k)-l_{0}v_{0})-k((l_{0}-k(t_{1}-t_{0}))-l_{0})$
$=2kl_{0}-kv_{0}(t_{1}-t_{0})-k^{2}(t_{1}-t_{0})$
$=k(2l_{0}-k(t_{1}-t_{0})-v_{0}(t_{1}-t_{0}))$
$=0,$
where we have used the fact that
$\begin{cases}l_{1}=l_{0}+\dot{l}_{0}(t_{1}-t_{0})\\ v_{1}=v_{0}-2\dot{l}_{0}\end{cases}$
Next in the time direction
$2\mathcal{T}^{-1}(\theta_{1}-\theta_{0})=\int_{t_{0}}^{t_{1}}\dfrac{ds}{l(s)^{2}}=\dfrac{t_{1}-t_{0}}{l_{0}l_{1}}\;.$
And meanwhile
$\dfrac{l_{0}l_{1}}{t_{1}-t_{0}}$ $=\dfrac{l_{0}v_{0}l_{1}}{l_{0}+l_{1}}$
$=\dfrac{l_{0}v_{0}(l_{0}+l_{1}-l_{0})}{l_{0}+l_{1}}$
$=l_{0}v_{0}-\dfrac{l_{0}(l_{1}-\dot{l}_{0}(t_{1}-t_{0}))}{t_{1}-t_{0}}$
$=\underbrace{l_{0}v_{0}+l_{0}\dot{l}_{0}}_{I_{0}/\mathcal{T}}-\dfrac{l_{0}l_{1}}{t_{1}-t_{0}}$
$=\dfrac{I_{0}}{\mathcal{T}}-\dfrac{l_{0}l_{1}}{t_{1}-t_{0}},$
which implies that
$\theta_{1}-\theta_{0}=\frac{1}{I_{0}}\;.$
2.1.2. The Adiabatic Normal Forms
Here, we present the adiabatic normal forms (c.f. Proposition 2.2), defined on the Poincaré sections $R_{0},R_{T}$, which capture the dramatic momentum change of the particle in roughly one period $\Delta t=2T$, due to the the singularities in the derivative of wall motions.
Starting with a high initial velocity, the particle experiences numerous collisions in a period. However, the adiabatic coordinates from Lemma 2.1 makes it possible to “sum up” these collisions until the particle foresees a singularity ahead of the next collision. We hereby perform another change of coordinates near the singularities and then complete the puzzle in the next proposition.
We define the “new time” variable $\tau$ as follows:
$\tau\buildrel\pmod{1}\over{:=}\begin{cases}I\theta&\hbox{in $R_{0}$}\\ I(\theta-\theta_{T})&\hbox{in $R_{T}$}\end{cases}$
where $\theta_{T}=\displaystyle\dfrac{1}{2\mathcal{T}}\int_{0}^{T}\dfrac{ds}{l(s)^{2}}=\dfrac{1}{2}$.
Proposition 2.2 (Adiabatic normal forms).
For all initial conditions $(\tau,I)$ with sufficiently large initial momentum $I$, the Poincaré maps of the particle are given by the following two adiabatic normal forms $P_{1}:R_{0}\to R_{T}$ and $P_{2}:R_{T}\to R_{0}$
(4) $P_{1}:=\begin{cases}\bar{\tau}=\tau-\dfrac{I}{2}\;\pmod{1}\\ \bar{I}=I+2Ak\mathcal{T}(2\bar{\tau}-1)\end{cases},$
(5) $P_{2}:=\begin{cases}\bar{\tau}=\tau-\dfrac{I}{2}\;\pmod{1}\\ \bar{I}=I-2Bk\mathcal{T}(2\bar{\tau}-1)\end{cases}.$
Proof.
We present the proof for the Poincaré map $P_{1}$; the proof for $P_{2}$ can be done in a similar fashion.
[…]
Figure 2. The Poincaré sections and the adiabatic normal forms
Let $(\theta_{0},I_{0})\in R_{0}$. We denote for ease $\tau=\theta_{0}I_{0},I=I_{0}$ and $(\bar{\tau},\bar{I})=P_{1}(\tau,I)$.
By Lemma 2.1, $(\theta_{n},I_{n})\in f^{-1}R_{0}$ with $n=[I_{0}(\theta_{T}-\theta_{0})]$. Then $\bar{\tau}=I_{n+1}(\theta_{n+1}-\theta_{T}),\bar{I}=I_{n+1}$.
Our goal is to derive formulas of $\bar{\tau},\bar{I}$ in terms of $\tau,I$.
For $t_{n+1}\in(T,2T)$, $l(t_{n+1})=A+k(t_{n+1}-T),v_{n+1}=v_{n}-2k$, so we have
(6) $\begin{cases}\bar{\tau}=\dfrac{I_{n+1}}{2\mathcal{T}}\displaystyle\int_{0}^{1}\dfrac{ds}{l(s)^{2}}=\dfrac{I_{n+1}}{2\mathcal{T}}\dfrac{t_{n+1}-T}{Al_{n+1}}\\ \bar{I}=\mathcal{T}(l_{n+1}v_{n+1}+l_{n+1}\dot{l}_{n+1})=\mathcal{T}l_{n+1}(v_{n+1}+k)=\mathcal{T}l_{n+1}(v_{n}-k)\end{cases},$
which implies that
(7) $\bar{\tau}=\dfrac{\mathcal{T}l_{n+1}(v_{n}+k)}{2\mathcal{T}}\dfrac{t_{n+1}-T}{Al_{n+1}}=\dfrac{1}{2A}(v_{n}-k)(t_{n+1}-T),$
which then suggests that we arrange computation in terms of $(v_{n}-k)$ and $(t_{n+1}-T)$.
We start with the formula for the momentum change.
We recall that
(8) $I_{n}=\mathcal{T}(l_{n}v_{n}+l_{n}\dot{l}_{n})=\mathcal{T}l_{n}(v_{n}-k).$
Then
$I_{n+1}$ $=\mathcal{T}l_{n+1}(v_{n}-k)$
$=\mathcal{T}(l_{n+1}-A+A-l_{n}+l_{n})(v_{n}-k)$
[…]
$=I_{n}+2Ak\mathcal{T}\bar{\tau}-k\mathcal{T}(T-t_{n})(v_{n}-k).$
Now let us simplify $(T-t_{n})(v_{n}-k)$.
$v_{n}(t_{n+1}-t_{n})$ $=l_{n}+l_{n+1}$
$\implies v_{n}(t_{n+1}-T+T-t_{n})$ $=(l_{n+1}-A)-(A-l_{n})+2A$
[…]
(9) $\implies(v_{n}-k)(T-t_{n})$ $=-2A(\bar{\tau}-1).$
Hence we continue
$I_{n+1}$ $=I_{n}+2Ak\mathcal{T}\bar{\tau}+2Ak\mathcal{T}(\bar{\tau}-1)$
$=I_{n}+2Ak\mathcal{T}(2\bar{\tau}-1)$
which implies, together with $I_{n}=I_{0}$ by Lemma 2.1, that
$\bar{I}=I+2Ak\mathcal{T}(2\bar{\tau}-1)\;.$
Next we derive the formula for the time change.
We note that in the $n^{th}$ collision right before the singularity at $t=T$
$I_{n}\theta_{n}$ $=I_{n}(\theta_{T}-(\theta_{T}-\theta_{n}))$
$=I_{n}\theta_{T}-\mathcal{T}(l_{n}v_{n}-kl_{n})\dfrac{1}{2\mathcal{T}}\int_{t_{n}}^{T}\dfrac{ds}{l(s)^{2}}$
$=\dfrac{1}{2}I_{0}-\dfrac{1}{2}l_{n}(v_{n}-k)\dfrac{T-t_{n}}{Al_{n}}$
$=\dfrac{1}{2}I_{0}-\dfrac{1}{2A}(v_{n}-k)(T-t_{n}).$
Then by (9) we have
$\bar{\tau}$ $=1-\dfrac{1}{2A}(v_{n}-k)(T-t_{n})$
$=1+I_{n}\theta_{n}-\dfrac{1}{2}I_{0}$
$=1+I_{0}\left(\theta_{0}+\dfrac{n}{I_{0}}\right)-\dfrac{1}{2}I_{0}$
$=I_{0}\theta_{0}-\dfrac{1}{2}I_{0}+n+1$
=$\buildrel\pmod{1}\over{=}\tau-\dfrac{1}{2}I_{0}\;.$
Remark 2.3.
We emphasize that the adiabatic coordinates and normal forms are only valid for large energies. In fact, the first equation in (1) is not true if $v_{0}$ is so small that a re-collision occurs, i.e. the moving wall chases and collides the particle before the particle collides at the fixed wall.
2.2. Skew Product Structure under Resonance
In this subsection we show that the complete Poincaré map $P:=P_{2}\circ P_{1}$ possesses invariant circles. If furthermore the parameters $A,B$ satisfy resonance conditions, then the parabolic dynamics on the invariant circles can be decomposed as a skew product over an interval exchange map with integer change on the fibers.
2.2.1. Invariant Circles
Since $P_{1},P_{2}$ from Proposition 2.2 are (piecewise) affine maps with constant coefficients, the eigendirection of the complete Poincaré map $P=P_{2}\circ P_{1}$ is always preserved. Now we verify that the translations in the affine maps are compatible with the dynamics so that we obtain invariant circles of $P$ on the $(\tau,I)$-phase cylinder.
We note that
$DP_{1}=\begin{pmatrix}1&-\dfrac{1}{2}\\ 4Ak\mathcal{T}&1-2Ak\mathcal{T}\end{pmatrix},\;DP_{2}=\begin{pmatrix}1&-\dfrac{1}{2}\\ -4Bk\mathcal{T}&1+2Bk\mathcal{T}\end{pmatrix}$
and consequently
$DP=DP_{2}\cdot DP_{1}=\begin{pmatrix}1-2Ak\mathcal{T}&Ak\mathcal{T}-1\\ 4(B-A)k\mathcal{T}&1+2Ak\mathcal{T}\end{pmatrix}$
where we have used the fact that $\mathcal{T}=\dfrac{T}{AB}$ and $k=\dfrac{B-A}{T}$.
We further note that $\det(DP)=1$, and $DP$ has an eigenvalue $\lambda=1$ with an eigenvector $\mathbf{v}=(A,2(A-B))$. We propose the following candidate as the family $\{\mathcal{C}_{D}\}_{D\in[0,1)}$ of invariant circles of the complete Poincaré map $P$:
(10) $\mathcal{C}_{D}:=\left\{(\tau,I)\,\big{|}\,\tau+\dfrac{A}{2(B-A)}=D\pmod{1}\right\}.$
We claim that $\mathcal{C}_{D}$ is invariant under the dynamics $P$.
Lemma 2.4 (Invariant circles).
For any $D\in[0,1)$, $\mathcal{C}_{D}$ as defined in (10) is invariant by the map $P$.
Proof.
Fix $D\in(0,1)$. Suppose $(\tau,I)\in\mathcal{C}_{D}$, i.e. $\tau+\dfrac{A}{2(B-A)}I=D\pmod{1}$. We denote $(\bar{\tau},\bar{I})=P_{1}(\tau,I)$ and $(\bar{\bar{\tau}},\bar{\bar{I}})=P_{2}(\bar{\tau},\bar{I})$.
By Proposition 2.2, we have
$\bar{\tau}$ $=\tau-\dfrac{I}{2}\mod 1$
$\bar{I}$ $=I+2Ak\mathcal{T}\left(2\left(\tau-\dfrac{I}{2}\mod 1\right)-1\right)$
$=I+4Ak\mathcal{T}\left(\tau-\dfrac{I}{2}\mod 1\right)-2Ak\mathcal{T}$
and then
$\bar{\bar{\tau}}$ $=\bar{\tau}-\dfrac{\bar{I}}{2}\mod 1$
$=\left(\tau-\dfrac{I}{2}\mod 1\right)-\dfrac{I}{2}-2Ak\mathcal{T}\left(\tau-\dfrac{I}{2}\mod 1\right)+Ak\mathcal{T}\mod 1$
$=\tau-I-2Ak\mathcal{T}\left(\tau-\dfrac{I}{2}\mod 1\right)+Ak\mathcal{T}\mod 1$
$\bar{\bar{I}}$ $=\bar{I}-2Bk\mathcal{T}(2\bar{\bar{\tau}}-1)$
$=I+4Ak\mathcal{T}\left(\tau-\dfrac{I}{2}\mod 1\right)-2Ak\mathcal{T}+2Bk\mathcal{T}$
$\quad-4Bk\mathcal{T}\left(\tau-I-2Ak\mathcal{T}\left(\tau-\dfrac{I}{2}\mod 1\right)+Ak\mathcal{T}\mod 1\right)\;.$
We verify that $(\bar{\bar{\tau}},\bar{\bar{I}})\in\mathcal{C}_{D}$.
$\bar{\bar{\tau}}+\dfrac{A}{2(B-A)}\bar{\bar{I}}$ $=_{1}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tau}-I-2Ak\mathcal{T}\left(\tau-\dfrac{I}{2}\mod 1\right)+Ak\mathcal{T}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\dfrac{A}{2(B-A)}I}$
$\quad+\underbrace{\dfrac{2A^{2}k\mathcal{T}}{B-A}}_{2A/B}\left(\tau-\dfrac{I}{2}\mod 1\right)-\underbrace{\dfrac{A^{2}k\mathcal{T}}{B-A}}_{A/B}+\underbrace{\dfrac{ABk\mathcal{T}}{B-A}}_{1}$
$\quad-\underbrace{\dfrac{2ABk\mathcal{T}}{B-A}}_{2}\left(\tau-I-2Ak\mathcal{T}\left(\tau-\dfrac{I}{2}\mod 1\right)+Ak\mathcal{T}\mod 1\right)$
$=_{1}D+\left(\tau-\dfrac{I}{2}\mod 1\right)\underbrace{\left(-2Ak\mathcal{T}+\dfrac{2A}{B}+4Ak\mathcal{T}\right)}_{2}$
$\quad+I-2\tau+\underbrace{Ak\mathcal{T}-\dfrac{A}{B}-2Ak\mathcal{T}+1}_{0}$
$=_{1}D+2\tau-I+I-2\tau$
$=_{1}D\;.$
2.2.2. Skew Product Structure
Now, we show that under resonance condition, the invariant circle $\mathcal{C}_{D}$ admits a skew product structure and the dynamics on $\mathcal{C}_{D}$ is equivalent to a finite extension of circle rotations with a piecewise constant cocycle.
Definition 2.5 (Classical resonance).
For $q\in\mathbb{N}$, we say the parameters $(A,B)$ are in (classical) $q$-resonance if $(B-A)/A=q$.
By our assumption $B>A$, $q\geq 1$. We say $(A,B)$ is in special resonance if $(B-A)/A=1$. The particular choice of parameters $(A,B,T)=(1/\sqrt{2},\sqrt{2},1)$ made by Ulam [Ulam61] and Zharnitsky [Zha98] corresponds to a special case of the special resonance.
We introduce some necessary notations before we state the main proposition of the section.
Definition 2.6 (A/typical circles).
We say an invariant circle $\mathcal{C}_{D}$ is typical if $D\neq m/q$ for any integer $m=0,\cdots,q-1$. Otherwise, we say $\mathcal{C}_{D}$ is atypical if $D=m/q$ for some integer $m=0,\cdots,q-1$.
We cut an invariant circle $\mathcal{C}_{D}$ into continuity components of $P_{1},P_{2}$ (c.f. Fig.LABEL:fig:contcomp).
Fix an integer $m=0,\cdots,q-1$.
For a typical circle $\mathcal{C}_{D}$ with $D\in\left(\dfrac{m}{q},\dfrac{m+1}{q}\right)$, we define for any $s=0,1,\cdots,q+1$
(11) $\mathcal{C}_{D}^{m,s}:=\mathcal{C}_{D}\cap\left\{\dfrac{q}{q+1}D-\dfrac{1}{q+1}(m+1-s)<\tau<\dfrac{q}{q+1}D-\dfrac{1}{q+1}(m-s)\right\}.$
We need to introduce the secondary cut on the starting or ending component, depending on whether $D$ lies in the first or the second half of $\big{(}m/q,(m+1)/q\big{)}$.
For $D\in\left(\dfrac{m}{q},\dfrac{m+\frac{1}{2}}{q}\right]$, we perform secondary cut on the ending component $\mathcal{C}_{D}^{m,q+1}$ at $\tau^{m,q+1}:=2-\frac{q}{q+1}(1-2D)-\frac{2}{q+1}(m+1)$
(12) $\mathcal{C}_{D}^{m,q+1-}:=\mathcal{C}_{D}^{m,q+1}\cap\{\tau<\tau^{m,q+1}\},\;\mathcal{C}_{D}^{m,q+1+}:=\mathcal{C}_{D}^{m,q+1}\cap\{\tau>\tau^{m,q+1}\}.$
For $D\in\left(\dfrac{m+\frac{1}{2}}{q},\dfrac{m+1}{q}\right)$, we perform secondary cut on the starting component $\mathcal{C}_{D}^{m,0}$ at $\tau^{m,0}:=1-\frac{q}{q+1}(1-2D)-\frac{2}{q+1}(m+1)$
(13) $\mathcal{C}_{D}^{m,0-}:=\mathcal{C}_{D}^{m,0}\cap\{\tau<\tau^{m,0}\},\;\mathcal{C}_{D}^{m,0+}:=\mathcal{C}_{D}^{m,0}\cap\{\tau>\tau^{m,0}\}.$